OpenStudy (anonymous):

Prove the following assertion: For all natural numbers n≥1, the expression [(2^{n}+1)(2^{n}-1)/3(2^{n}+(-1)^{n})\] is a positive interger.

OpenStudy (anonymous):

the question looks like this: Prove the following assertion:\[(2^{n}+1)(2^{n}-1)/3(2^{n}+(-1)^{n})\]is a positive integer

OpenStudy (anonymous):

(2^n +1)/3 , when n is odd which is positive for all valus of n.. again (2^n - 1)/3, when n is even , which is always positive if n>1.

OpenStudy (anonymous):

thank you! Are there any particular steps to follow to make it a little clearer?

OpenStudy (anonymous):

step 1. when n is odd..then (-1)^n=-1.. plug the value, u get (2^n +1)(2^n -1)/3(2^n -1) = (2^n +1)/3. step 2, when n is even ...then (-1)^n= 1. similarly plug this value...you will get (2^n -1)/3...... you got it???

OpenStudy (anonymous):

I got it now, thanks :)